OpenStudy (akashdeepdeb):

We are given four equations and we have to find n here: \[4n = m^2 + k^2 + m + k\] \[8056 = m^2 - k^2 + m - k\] \[n - 2014 = \frac{k(k+1)}{2}\] \[n + 2014 = \frac{m(m+1)}{2}\] btw, I derived the first 2 equations from the bottom 2 equations.

OpenStudy (akashdeepdeb):

@ganeshie8 @AravindG

OpenStudy (akashdeepdeb):

@robtobey @SolomonZelman

OpenStudy (solomonzelman):

What were the original equations?

OpenStudy (akashdeepdeb):

The bottom two.

OpenStudy (anonymous):

its solvable :)

OpenStudy (anonymous):

i got 2n=1 n=1/2

OpenStudy (akashdeepdeb):

n has to be a natural number.

OpenStudy (akashdeepdeb):

:/

OpenStudy (anonymous):

wait :3

OpenStudy (akashdeepdeb):

I got 2 x 2 x 2 x 19 x 53 = (m+k+1)(m-k) Now (m+k+1) is always greater than (m-k) We CAN use this to find 'n' but it'll take a heck lot of time!

OpenStudy (anonymous):

sorrt idk lol @ganeshie8 wanna try ?

OpenStudy (akashdeepdeb):

@agent0smith @nincompoop

OpenStudy (solomonzelman):

Idk, but to me it looks like you have to think of isolating one of the variables using "what 2014 is equal to", but other than that it is 2 equations with 3 variables. Also knowing that it has to be positive integer... too much thinking for me, sorry -;(

OpenStudy (anonymous):

its not like that lol , but i know nothing abt k,m how could i continue ?

OpenStudy (anonymous):

whats the whol question cuz i dnt read the horoscope lolz

OpenStudy (anonymous):

Now (m+k+1) is always greater than (m-k) , only when m,k possitive

ganeshie8 (ganeshie8):

yes also : 2 x 2 x 2 x 19 x 53 = (m+k+1)(m-k) you have only 16 cases to consider

ganeshie8 (ganeshie8):

so why do u think it gona take lot of time ? and an observation - right hand side of both equations is a triangular number

OpenStudy (akashdeepdeb):

But I'd have to solve for m and k, 16 times, and i may get n only few times. m and k and n are all greater than zero.

OpenStudy (anonymous):

ohh lol so n-2014 n+2014 should apply 

ganeshie8 (ganeshie8):

Also notice that if m-k is even, then m+k+1 has to be odd and vice versa

OpenStudy (akashdeepdeb):

why?

OpenStudy (akashdeepdeb):

oh got it ok.

ganeshie8 (ganeshie8):

so that reduces our work by more than a half :)

OpenStudy (akashdeepdeb):

Now how many cases are going to be there?

ganeshie8 (ganeshie8):

the even prime powers : 2^3 must all lie in either m-k or m+k+1

ganeshie8 (ganeshie8):

m-k and m+k+1 cannot share 2's

OpenStudy (akashdeepdeb):

Oh yeah! So 8 * 19 * 53 are the options now right?

ganeshie8 (ganeshie8):

Looks good ! 8 cases to consider now

OpenStudy (akashdeepdeb):

3 cases right? ;____; 8*19 * 53 19*53 * 8 8*53 * 19

OpenStudy (akashdeepdeb):

@ganeshie8 ?

ganeshie8 (ganeshie8):

m-k = {1, 8,19, 53, 8*19, 19*53, 53*8, 8*19*53}

ganeshie8 (ganeshie8):

dont be so much relieved... :P there are 8 cases to work^^

OpenStudy (akashdeepdeb):

Oh yeah, how could i forget the '1'. Yeah, not relieved at all. ;__;

ganeshie8 (ganeshie8):

i dont see any easy ways out... u need to test all 8 cases i think... goodluck !

OpenStudy (akashdeepdeb):

Thanks but, you helped me narrow it down to 8 cases. T___T

ganeshie8 (ganeshie8):

use that to verify/reverse engineer ^

OpenStudy (akashdeepdeb):

m and k both always have to positive right? Because they form, the triangular numbers?

ganeshie8 (ganeshie8):

I think so, i never heard of negative triangular numbers... @BSwan do u ?

OpenStudy (anonymous):

nope